OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 2x - 3y = 4 x + 4y = -9

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

just use exactly the same rules as we showed you for the other questions

OpenStudy (anonymous):

x,y=2.2,-2.8

OpenStudy (anonymous):

x = -9 - 4y 2(-9-4y)-3y = 4 -18 - 8y - 3y = 4 -18 - 11y = 4 -11y = 22 y = -2 x = -1